OpenStudy (anonymous):

[URGENT= will give BADGE] I need help finding the wronskian of these four functions...

OpenStudy (anonymous):

\[y= 1, y = t, y= \cos (t), y= \sin (t)\]

OpenStudy (anonymous):

@nader1

OpenStudy (anonymous):

isnt it 4x5, since the funtion goes on top?

OpenStudy (anonymous):

or does the equation given with matter at all?

OpenStudy (anonymous):

like \[y ^{4} + y ^{2} = 0\]

OpenStudy (anonymous):

i guess that part is confusing me...

OpenStudy (anonymous):

ok im given that equation above and give 4 solutions that I checked are solutions and now I have to find their Wronskian. Can you help?

OpenStudy (anonymous):

@Hoa

OpenStudy (anonymous):

those are derivatives by the way.

OpenStudy (anonymous):

\[y \prime \prime \prime \prime + y \prime \prime = 0\]

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

can u just answer me this:

OpenStudy (anonymous):

can u please write out the 2nd part of the Wronskian only

OpenStudy (anonymous):

i already know the first stage...but am messing up somewhere on the 2nd stage

OpenStudy (anonymous):

when it's down to the 3x3 matrices

OpenStudy (anonymous):

only the first term matter in the 2nd stage, the rest are zeros.

OpenStudy (anonymous):

so can u help me?

OpenStudy (anonymous):

im getting the wrong sign at the end...

OpenStudy (anonymous):

\[\cos (t) - \sin (t) \]

OpenStudy (anonymous):

The answer should be 1.

OpenStudy (anonymous):

i meant \[\cos ^2 (t) - \sin ^2 (t)\]

OpenStudy (nincompoop):

sorry, my computer wasn't loading properly

OpenStudy (nincompoop):

pretty @FibonacciChick666 got this though. she's good ;)

OpenStudy (fibonaccichick666):

haha thanks @nincompoop , so i'm just figuring out how to write this give me a sec.

OpenStudy (fibonaccichick666):

ok so I interpreted this as a parameterization wrt t. this gives us a 3x3 matrix for calculating the wronskian. as shown below \[\left[\begin{matrix}t & \cos t & sint \\1 & -sint & \cos t\\ 0 & -\cos t & -sint \end{matrix}\right]\]

OpenStudy (fibonaccichick666):

does this help? and do you see the answer now?